# MonoPCC: Photometric-invariant Cycle Constraint for Monocular Depth Estimation of Endoscopic Images

### v1

@@ 5.2.1 Effectiveness of Three Components
Table 5 lists the comparison results as well as the backbone MonoViT using none of these techniques. As can be seen, the variant using STM+EMA strategy exceeds that using AFM only with a reduction of Abs Rel and Sq Rel by 7.14% and 13.59%, respectively ($p=0.019$, $p=0.014$). It is worth noting that compared to AFM, our method requires no extra model to learn, exhibiting greater usability and scalability.
@@ 6 Conclusion
Self-supervised monocular depth estimation is challenging for endoscopic scenes due to the severe negative impact of brightness fluctuations on the photometric constraint. In this paper, we propose a cycle-form warping to naturally overcome the brightness inconsistency of endoscopic images, and develop a MonoPCC for robust monocular depth estimation by using a re-designed photometric-invariant cycle constraint. To make the cycle-form warping effective in the photometric constraint, MonoPCC is equipped with two enabling techniques, i.e., structure transplant module (STM) and exponential moving average (EMA) strategy. STM alleviates image detail degradation to validate the backward warping, which uses the result of forward warping as input. EMA bridges the learning of network weights in the forward and backward warping, and stabilizes the intermediate warped image to ensure an effective convergence. The comprehensive and extensive comparisons with $8$ state-of-the-arts on three public datasets, i.e., SCARED, SimCol3D, and SERV-CT, demonstrate that MonoPCC achieves a superior performance by decreasing the absolute relative error by at least $7.27\%$. Additionally, two ablation studies are conducted to confirm the effectiveness of three developed modules and the advancement of MonoPCC over other similar techniques against brightness fluctuations. In the future work, we plan to extend MonoPCC into the multi-frame based monocular depth estimation.

### v3

@@ 5.2 Ablation Study
5.2.1 Effectiveness of Three Components
We develop five variants of MonoPCC by disabling the three components, EMA and/or $\mathcal{L}_{pcp}$, and STM. Note that, the warping should be cycle-form to utilize the components, and if we disable EMA, only $\theta$ in the backward warping path will be updated, and $\theta_{EMA}$ in the forward warping remains unchanged. Table 4 and Fig. 7 give the comparison results between the variants and complete MonoPCC. We also include the backbone MonoViT in the first row of Table 4 as the baseline, which adopts the non-cycle warping.
From the comparison results in Table 4 and visualization cases in Fig. 7, three key observations can be made:
(1) By comparing the first two rows of Table 4 and Fig. 7 (a)-(b), the inferior performance of the variant using none of the components indicates that the direct usage of cycle warping is not enough due to the image blurring and unstable gradient propagation mentioned in Sec. 3.2.
(2) As can be seen from the second to fourth rows of Table 4 and Fig. 7 (b)-(d), using EMA and $\mathcal{L}_{pcp}$ solely brings no improvements, which still creates much unsatisfactory estimation in the red dotted boxes of Fig. 7 (c). However, adding STM can significantly reduce Sq Rel and RMSE by 16.11% and 7.50%, respectively ($p<0.003$). Thus, STM is the irreplaceable component for successfully training using the cycle-form warping.
(3) Comparison between the last four rows of Table 4 and Fig. 7 (d)-(g) reveals the effectiveness of EMA and $\mathcal{L}_{pcp}$. Specifically, with STM enabled, both EMA and $\mathcal{L}_{pcp}$ bring remarkable reduction of Abs Rel by 8.08% ($p=0.012$, $p=0.022$). Meanwhile, EMA and $\mathcal{L}_{pcp}$ are not mutually excluded, and the combination reduces Abs Rel by 14.14% ($p=0.001$). As illustrated in Fig. 7 (g), there exist more blue-tuned pixels in the red dotted regions of interest (i.e., either light or dark areas).
@@ 5.2.1 Effectiveness of Three Components
Table 5 lists the comparison results as well as the backbone MonoViT using none of these techniques. As can be seen, the variant using STM+EMA strategy exceeds that using AFM only with a reduction of Abs Rel and Sq Rel by 7.14% and 13.59%, respectively ($p=0.019$, $p=0.014$). Fig. 8 displays two visualization examples in challenging cases, e.g., the boundary and flat area of tissues. As shown in Fig. 8, all techniques for addressing the brightness fluctuation can help to gain varying degrees of improvements. Furthermore, we observe that our strategy performs better than both ABT and AFM. It is worth noting that compared to AFM, our method requires no extra model to learn, exhibiting greater usability and scalability.
@@ 5.3 Robustness to Severe Brightness Inconsistency
Figure 8: The Abs Rel error maps of MonoPCC and other similar modules against photometric inconsistency. (a)-(d) correspond to the $1^{st}$ to the $4^{th}$ rows in Table 5. Opencv Jet Colormap is used for visualization.
@@
Table 6: Quantitative comparison results on KITTI. The best results are marked in bold and the second best ones are underlined.
Year Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$ $\delta$uparrow
Monodepth2 2019 0.115 0.903 4.863 0.193 0.877
FeatDepth 2020 0.104 0.729 4.481 0.179 0.893
HR-Depth 2021 0.109 0.792 4.632 0.185 0.884
DIFFNet 2021 0.102 0.749 4.445 0.179 0.897
Endo-SfMLearner 2021 0.100 0.722 4.457 0.178 0.898
AF-SfMLearner 2022 0.113 1.110 4.949 0.189 0.883
MonoViT 2022 0.099 0.708 4.372 0.175 0.900
Lite-Mono 2023 0.101 0.729 4.454 0.178 0.897
MonoPCC(Ours) 0.098 0.677 4.318 0.173 0.900
5.4 Results on KITTI Benchmark
Besides the demonstration on the medical scenarios, i.e., endoscopic images, we in this subsection verify the competitiveness of MonoPCC on the popular natural-scence benchmark, i.e., KITTI [7]. KITTI is an outdoor dataset, which roughly obeys brightness consistency assumption due to constant illumination condition, i.e., the sunshine. However, some factors like specular reflection on the car body also create some negative impacts on the ideal photometric supervision.
Table 6 provides the comparison results with the previously mentioned SOTAs. Most of them have reported their evaluation results on KITTI, so we directly copy their results. For the two methods, i.e., Endo-SfMLearner and AF-SfMLearner, which have no official results on KITTI, we use their released codes and default settings to train two models on KITTI.
As indicated in Table 6, MonoPCC and MonoViT are the only two methods whose Abs Rel values are less than $0.1$. Moreover, MonoPCC outperforms MonoViT in terms of all metrics, especially with a relative decrease of 4.38% in terms of Sq Rel error ($p<0.001$).
The two endoscopy-tailored methods, i.e., Endo-SfMLearner and AF-SfMLearner, are the second or the third best on the four endoscopic datasets, but unexpectedly slip several places on KITTI. We believe the potential reason is that the two methods are overly focusing on the brightness adjustment, and try to change the image appearance no matter the brightness is inconsistent or not. Since KITTI has no that significant brightness fluctuations, such forcible brightness adjustment could induce errors more or less. In comparison, our MonoPCC is essentially making the brightness patterns inherit across the source and target images, and thus the strength of brightness adjustment can be naturally compatible with the true situation of brightness fluctuations. Therefore, MonoPCC shows the consistent competitiveness in both medical and natural scenes.
@@ 6 Conclusion
Self-supervised monocular depth estimation is challenging for endoscopic scenes due to the severe negative impact of brightness fluctuations on the photometric constraint. In this paper, we propose a cycle-form warping to naturally overcome the brightness inconsistency of endoscopic images, and develop a MonoPCC for robust monocular depth estimation by using a re-designed photometric-invariant cycle constraint. To make the cycle-form warping effective in the photometric constraint, MonoPCC is equipped with two enabling techniques, i.e., structure transplant module (STM) and exponential moving average (EMA) strategy. STM alleviates image detail degradation to validate the backward warping, which uses the result of forward warping as input. EMA bridges the learning of network weights in the forward and backward warping, and stabilizes the intermediate warped image to ensure an effective convergence. The comprehensive and extensive comparisons with $8$ state-of-the-arts on five public datasets, i.e., SCARED, SimCol3D, SERV-CT, Hamlyn, and KITTI, demonstrate that MonoPCC achieves a superior performance by decreasing the absolute relative error by at least 7.27%, 9.38%, 9.90% and 3.17% on the four endoscopic datasets, respectively, and shows the competitiveness even for the natural scenario. Additionally, two ablation studies are conducted to confirm the effectiveness of three developed modules and the advancement of MonoPCC over other similar techniques against brightness fluctuations. However, the current pipeline only leverages a single frame to infer the depth map, which is likely to lack context information and geometric consistency. In the future work, we plan to extend MonoPCC into the multi-frame based monocular depth estimation.
Acknowledgments
This work was supported in part by National Key R&D Program of China (Grant No. 2023YFC2414900), Fundamental Research Funds for the Central Universities (2021XXJS033), Research grants from United Imaging Healthcare Inc.